OpenStudy (anonymous):

A quadratic equation is shown below: x2 + 5x + 4 = 0 Part A: Describe the solution(s) to the equation by just determining radicand. Show your work. (3 points) Part B: Solve 4x2 -12x + 5 = 0 using an appropriate method. Show the steps of your work, and explain why you chose the method used. (4 points) Part C: Solve 2x^2 -10x + 3 = 0 by using a method different from the one you used in Part B. Show the steps of your work. (3 points)

OpenStudy (anonymous):

I have Part A and B but i need C, For B I used the formula -b+-b^2-4ac/2a. Now I just need to factor, but how do you do that?

OpenStudy (anonymous):

you could use completing the square.

OpenStudy (anonymous):

2x^2 -10x + 3 = 0 2(x^2 -5x + (b/2)^2) = -3 b is -5 2(x^2 -5x + 25/4) = -3 + 25/2 you add 25/2 because you add 25/2 to the other side ( it was multiplied by 2 so its 25/2 not 25/4) 2(x-5/2)^2 = 19/2 divide by 2 (x-5)^2 = 19/4

OpenStudy (anonymous):

and you can do the rest... sorry i have to go. but if you dont understand i will check on it soon

OpenStudy (anonymous):

okay since i came back before send anything i will finish the problem. square root both side. x-5= +-sqrt(19/4) x= \[5 \pm \frac{ \sqrt{19} }{ 2 }\]